Please help with the following question:
y-3=3(x+1)
A.) what is the equation in standard form of a parallel line that passes through (0,-3).
B.) On the parallel line, find the ordered pair x=2.
Thank you greatly!

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Parallel to what?

Answer by stanbon(75887)

y-3=3(x+1)
y = 3x+6
slope = 3
y-int = 6
--------------------
A.) what is the equation in standard form of a parallel line that passes through (0,-3).
----
Form:: y = mx + b
Solve for "m" and for "b"::
m = 3 so the lines will be parallel
y-int = f(0) = -3 so b = -3
-----
Ans: y = 3x-3
-----------------
B.) On the parallel line, find the ordered pair x=2.
f(2) = 3*2-3 = 3
------------------
Cheers,
Stan H.
---------